OpenStudy (anonymous):

cos A = 5/13 Find the following: cos 2A, cos 4A, tan 2A, sin (A/2)

OpenStudy (anonymous):



OpenStudy (anonymous):

cos2A=2(cosA)^2-1 cosA=5/13

OpenStudy (anonymous):

cos2A=2*(5/13)^2-1